OpenStudy (anonymous):

What is the reciprocal of a mixed fraction? for example 1 1/2

OpenStudy (wwe123):

3/2 2/3

OpenStudy (anonymous):

1 1/2 is not right it should be 1 1 or 1/2 2/1 their multiplication will give always 1 eg 3/2 * 2/3 = 1 or 1/2 * 2 = 1 123/321 * 321/123 = 1

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20090423224408AAitZ5f check this

OpenStudy (anonymous):

thanks Vogelfrei